OpenStudy (anonymous):

Simplify. 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2 2 square root of 8 8 square root of 2 8 square root of 6 6 square root of 8

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (darkbluechocobo):

8 root of 2\[4\sqrt2+7\sqrt2-3\sqrt2\]

OpenStudy (darkbluechocobo):

just simple addition

OpenStudy (darkbluechocobo):

since they have the same sqrt

OpenStudy (anonymous):

Okay so how do i do that

OpenStudy (darkbluechocobo):

4+7-3

OpenStudy (anonymous):

so 8?

OpenStudy (darkbluechocobo):

yes \[8\sqrt2\]

OpenStudy (anonymous):

could you help me with a couple more?

OpenStudy (darkbluechocobo):

Sure

OpenStudy (anonymous):

Simplify.7 square root of 3 end root minus 4 square root of 6 end root plus square root of 48 end root minus square root of 54 11 square root of 6 end root minus 7 square root of 12 11 square root of 3 end root minus 7 square root of 6 negative 3 square root of 9 4 square root of 9

OpenStudy (anonymous):

Simplify. square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses 14 15 square root of 2 5 square root of 2 end root minus 4 square root of 10 None of the above

OpenStudy (darkbluechocobo):

2nd one is b

OpenStudy (darkbluechocobo):

7√3 - 4√6 + 4√3 - 3√6 = (7√3 + 4√3) + (- 4√6 - 3√6) 11√3 - 7√6

OpenStudy (anonymous):

i don't understant

OpenStudy (anonymous):

*understand

OpenStudy (darkbluechocobo):

basically \[7\sqrt3 - 4\sqrt6 +\sqrt(16*3)-\sqrt(9*6)=7\sqrt3 - 4\sqrt6 +\sqrt16 *\sqrt3 - \sqrt9 *\sqrt 6\]

OpenStudy (darkbluechocobo):

just have to break down the ones that can still be simplified

OpenStudy (anonymous):

how do i do that?

OpenStudy (darkbluechocobo):

so you saw that you had sqrt54 the sqrt of 54 would be simplified by using sqrt9*6 because that equals that. then the sqrt9 can be alone because the sqrt of 9 is three however with 6 you cannot simplifie it any more

OpenStudy (anonymous):

okay so the answer would be c?

OpenStudy (anonymous):

@DarkBlueChocobo

OpenStudy (anonymous):

@gabgurl

OpenStudy (darkbluechocobo):

B because it all simplifies to 7√3 - 4√6 + 4√3 - 3√6 = (7√3 + 4√3) + (- 4√6 - 3√6) 11√3 - 7√6 so \[11\sqrt3 -7\sqrt6\] so b is your answer

OpenStudy (anonymous):

oh okay thank you